# Beyond Rate Coding: Surrogate Gradients Enable Spike Timing Learning in Spiking Neural Networks

### v1

@@ 3.1 Construction of Timing-Normalized SHD and SSC Datasets
Spiking neural network (SNN) benchmarks such as SHD and SSC exhibit complex spatiotemporal spike patterns (Cramer et al., 2020). However, their raw versions suffer from pronounced spike-count biases: some channels fire far more often than others, and overall activity levels vary significantly between classes. Consequently, even simple multilayer perceptrons trained only on per-sample spike counts can achieve nontrivial accuracy (see fig. 5), making it hard to determine whether SNNs truly leverage precise timing or simply exploit rate-based cues.
Figure 3: Example spike rasters of SHD (top) and SSC (bottom) samples across the Whole $\rightarrow$ Part $\rightarrow$ Norm normalization stages. Neuron selection and spike count normalization progressively remove rate-based information while retaining temporal structure.
@@
Stage 1: Neuron and Sample Filtering for Spike Count Normalization (Whole $\rightarrow$ Part). To eliminate rate-based information while preserving temporal structure, we aim to normalize the spike count for each neuron across all samples. However, some neurons emit zero spikes in certain samples, making exact normalization infeasible.
@@
Neurons with $c_{i}^{\min}<\theta$ (we use $\theta=2$) are initially marked for exclusion. However, in many cases, a neuron’s low $c_{i}^{\min}$ is caused by a small number of inactive samples. To retain more neurons, we apply an additional sample filtering step: if a neuron’s low activity affects fewer than a threshold fraction of samples, we remove those samples instead of the neuron. Formally, for each neuron $i$, we define the set of affected samples:
@@
if |$\text{if }\frac{|\mathcal{M}_{i}|}{M}<\epsilon,\text{ then exclude all samplesin }\mathcal{M}_{i}.$ (8)
@@
After filtering, we retain only neurons that spike at least $\theta$ times in every remaining sample, forming the Part variant ($N^{\prime}\ll N$). To ensure a fair comparison across classes, we randomly downsample each class to match the size of the smallest one. This balancing step improves the interpretability of accuracy metrics in the subsequent experiments.
Stage 2: Min-Count Spike Normalization (Part $\rightarrow$ Norm). To fully eliminate spike count information while preserving spike timing, we subsample each spike train to retain exactly $c_{i}^{\prime\min}$ spikes per neuron $i$ in every sample, where $c_{i}^{\prime\min}$ is the minimum spike count across the filtered samples defined in Stage 1. Specifically, we randomly select $c_{i}^{\prime\min}$ spike times from each neuron’s spike train in each sample, ensuring fixed per-neuron spike counts across the dataset:
$x_{m,i,t}^{\text{norm}}=\texttt{Subsample}\big{(}\{t\mid x_{m,i,t}=1\},\,c_{i}^{\prime\min}\big{)}.$ (9)
@@
This process removes all rate-based information while retaining precise spike timing within each retained spike. However, it may distort global temporal patterns such as inter-spike intervals (ISIs) or synchrony between neurons, depending on which spikes are retained. The resulting Norm dataset preserves the shape of the Part variant and serves as a strict test of spike timing-based learning.
@@
3.2 Surrogate GD Enables SNNs to Exploit Spike Timing in Real Datasets
In this work, all experiments are conducted within the SLAYER framework (Shrestha and Orchard, 2018), comparing a baseline feedforward SNN trained with surrogate gradient descent (SGD) against an enhanced variant incorporating learnable axonal delays (SGD-delay) (Sun et al., 2023).
@@ 3.2 Surrogate GD Enables SNNs to Exploit Spike Timing in Real Datasets
Figure 4: Comparison of SNN architectures trained with SGD and SGD-delay. The top row shows a model with no delay components. The bottom row inserts learnable axonal delays between layers.
The two architectures are illustrated in fig. 4. Both models employ a two-layer SNN, each layer comprising 128 neurons, and are trained using the Spikemax loss function (Shrestha et al., 2022). In the SGD-delay model, axonal delays are left unbounded, allowing them to adapt freely to the data during training.
We compare SGD and SGD-delay training on the SHD and SSC datasets to evaluate their ability to extract temporal information from spike trains. Both methods are applied to three dataset variants (whole, part, norm), where the norm variant eliminates spike count differences across samples, isolating spike timing as the only useful cue.
As shown in fig. 5, both SGD and SGD-delay trained SNNs outperform the MLP baseline on the norm variant (red dashed lines), demonstrating their ability to extract information from spike timing alone. This confirms that surrogate gradient based-learning is not limited to learn rate-based features but can indeed learn from precise spike timing, even without explicit delay modeling.
@@
Figure 5: Test accuracy under spike timing perturbation $f$ on SHD (left) and SSC (right) datasets. Blue: whole; green: part; red: norm. For the norm variant, spike count information is fully removed, so MLP accuracy corresponds to chance level (5% for SHD, 2.8% for SSC).
Notably, SGD-delay consistently achieves higher accuracy than SGD across both datasets. For instance, on SHD-norm at $f=0$, SGD achieves 23.32% while SGD-delay reaches 47.86%; on SSC-norm, the scores are 14.69% and 32.60% respectively. These norm results serve as spike timing-only baselines and highlight the benefit of delay learning. The improvement reflects that learnable delay modules enable more accurate temporal credit assignment and introduce a stronger inductive bias toward spike timing—while incurring only minimal parameter increase: the SGD model has $mn$ parameters, while SGD-delay adds just $n$ additional delay terms, where $m$ and $n$ are the input and hidden dimensions.
@@
Across both SHD and SSC datasets, we observe a common trend in SGD-delay’s performance: as $f$ increases, the degradation in accuracy accelerates sharply—suggesting strong reliance on high-fidelity spike timing. Around $f{=}0.8$, the curves rapidly drop to the corresponding MLP baseline, after which the decay flattens out. This saturation behavior indicates that once spike timing is too corrupted to be useful, SGD-delay transitions to a rate-like regime, converging to MLP performance (chance level in the case of the norm variant).
4 SNNs Trained via Surrogate GD Are Sensitive to Temporal Structure and Reversal
4.1 Temporal Perturbations Induce Distinct but Predictable Degradation
To further evaluate the temporal sensitivity of surrogate gradient-trained SNNs, we move beyond the random spike replacement used in earlier sections (see eq. 2), and systematically design three biologically inspired temporal distortions: Per-spike and per-neuron jitter, and per-spike deletion. These perturbations target distinct aspects of spike-based encoding and allow us to assess how each type of timing disruption affects classification performance.
(a) Per-Spike Jitter. This perturbation mimics synaptic noise by adding zero-mean Gaussian noise to individual spike times, disrupting both inter-spike intervals (ISIs) and cross-neuron synchrony while preserving spike counts:
[…]
where $\delta_{i,j}\sim\mathcal{N}(0,\sigma^{2})$ and $\mathrm{clip}(\cdot)$ ensures spikes remain within bounds.
(b) Per-Neuron Jitter. We apply a constant Gaussian time offset to all spikes emitted by a neuron, preserving intra-neuron ISI structure while disrupting precise spike-time alignment across neurons:
[…]
with $\Delta_{i}\sim\mathcal{N}(0,\sigma_{N}^{2})$. This perturbation tests the model’s reliance on cross-neuron synchrony for decoding.
(c) Per-Spike Deletion. Each spike is independently deleted with probability $p_{d}$:
$\mathbb{P}[\text{delete }t_{i,j}]=p_{d}.$ (12)
This distortion degrades all timing-based representations by reducing both the amount and precision of temporal cues, including ISIs, synchrony, and rate. It serves as a strong test of model robustness under severe information loss.
@@ 4.1 Temporal Perturbations Induce Distinct but Predictable Degradation
Figure 6: SNN performance under temporal perturbations: (a) Per-spike jitter, (b) per-neuron jitter, (c) per-spike deletion. Top: spike rasters of one specific SHD-norm sample before (black) and after (red) perturbation. Bottom: test accuracy across three dataset variants (Whole, Part, Norm), using SGD with and without delay learning.
We evaluate model robustness under these perturbations on all three SHD variants: whole, part, and norm (see fig. 6). First, we compare per-spike versus per-neuron jitter at high noise levels ($\sigma=\sigma_{N}=25$ ms). Despite using the same Gaussian perturbation principle, the difference in granularity yields divergent results: in fig. 6(a), the model under per-spike jitter still achieves over 10% higher accuracy than the MLP baseline, whereas in fig. 6(b), per-neuron jitter degrades performance to MLP level. This suggests that per-spike jitter retains more useful temporal information. The key distinction lies in perturbation scope: per-spike jitter affects each spike individually, allowing some timing cues to survive, while per-neuron jitter shifts all spikes uniformly, eliminating relative spike timing across neurons and impairing cross-neuron synchrony more severely.
Although per-neuron jitter preserves intra-neuron ISIs by design, its performance across all dataset variants (whole, part, norm) is consistently worse than per-spike jitter, indicating that ISIs alone are not the dominant temporal feature driving classification. Instead, fine-grained inter-neuron timing appears more critical for model performance.
In contrast, fig. 6(c) shows that per-spike deletion disrupts both spike timing and count. As deletion probability increases, the performance of models trained on whole and part datasets drops below the MLP baseline, indicating that critical information—including spike redundancy and count—is lost. For the norm dataset, which has already been count-normalized, deletion leads to convergence near the chance level.
Overall, surrogate-trained SNNs show consistent degradation trends under timing perturbations, but their vulnerability depends strongly on the type of distortion. Fine-grained jitter preserves more information than coarse neuron-level shifts, while deletion, by directly removing spikes, proves the most destructive.
4.2 Vulnerable to Time Reversal Like Humans
@@
Humans are highly sensitive to the causal structure of sensory input—e.g., speech played backward becomes unintelligible (Saberi and Perrott, 1999). To examine whether SNNs exhibit similar temporal sensitivity, we conduct a time reversal experiment by reversing spike trains within their active time windows.
@@
$[t_{\mathrm{start}},\ t_{\mathrm{end}}]=\left[\min\{t\ |\ x_{i,t}=1\},\ \max\{t\ |\ x_{i,t}=1\}\right],$ (13)
@@
Figure 7: Test accuracy under spike timing perturbation $f$ for original and time-reversed SHD variants. (a) Spike raster of a single sample before and after temporal inversion. (b) Results for models trained with SGD. (c) Results for models trained with SGD-delay. Each curve corresponds to one of the whole, part, or norm datasets in both original and inverted forms.
We evaluate pretrained surrogate gradient models on time-reversed SHD variants to assess their sensitivity to reversed temporal structure. Without retraining, we test zero-shot generalization to time-inverted inputs by applying the transformation in eq. 14. Additionally, we introduce varying levels of random perturbation $f$ to examine whether noise can mitigate the impact of reversal.
Figure 7 reveals both reversal-induced vulnerability and reversal-invariant robustness in surrogate-trained SNNs. For the whole-inv dataset, accuracy is lowest at $f{=}0$ in both fig. 7(b) and (c), where precise yet reversed spike timing strongly misleads the model—especially for SGD-delay, which learns richer temporal dependencies. As $f$ increases and spike timings become progressively randomized, this misleading structure is weakened, and accuracy gradually recovers. This suggests that temporal noise suppresses harmful patterns and shifts the model’s reliance from spike timing to spike count.
In contrast, for the norm-inv datasets, the SGD-trained SNNs in both panels achieve accuracy above the MLP baseline at $f=0$. This suggests that SGD enables the models to learn spike timing features—such as inter-spike intervals (ISIs) or cross-neuron synchrony—that remain informative even when the temporal order of the input is reversed.
Interestingly, for both the SGD and SGD-delay trained models, the inverse variants exhibit a progressive trend relative to the MLP baseline: whole-inv performs below MLP level, part-inv remains close to MLP, and norm-inv achieves above-MLP accuracy. This monotonic shift suggests that the whole dataset contains a substantial amount of spike timing information that is sensitive to time reversal, which misleads the models when temporal order is inverted. As we move from whole to part then to norm, many of these vulnerable temporal features are progressively filtered out by construction.
These findings demonstrate that SNNs trained via surrogate gradient descent exhibit human-like sensitivity to temporal causality: they perform well when the input follows natural time order but misfire under reversal—mirroring the well-known phenomenon that time-reversed sensory inputs (e.g., speech) are often unintelligible to humans.
@@ 5 Discussion
In particular, we looked at the Spiking Heidelberg Digits and Spiking Speech Commands datasets (SHD, SSC). These have become widely used benchmarks in the field of spiking neural networks and neuromorphic computing. We find that these datasets do contain substantial and exploitable temporal information, but that it is also possible to perform surprisingly accurate classification using spike counts only, or following various types of perturbation of the spike timing. As a consequence, this dataset may not be ideal if we wish to probe the extent to which various spike-based models can exploit temporal information. We built a derived dataset (SHD-norm) which, by construction, eliminates all spike count information. (Note that although the only information remaining in this dataset is temporal, it is likely that a significant amount of temporal information is also removed by the transformations applied.) We have made a public release of this dataset to facilitate further research on timing-based computation in spiking neural networks.
@@
All code is available at the GitHub repository: https://github.com/neural-reckoning/temporal-shd. The datasets are archived on Zenodo(Yu et al., 2025) (https://doi.org/10.5281/zenodo.16153275).

### v2

@@ 3.2 Construction of Timing-Normalized SHD and SSC Datasets
While SHD and SSC exhibit complex spatiotemporal spike patterns (Cramer et al., 2020), spike count vectors vary significantly between classes and consequently, even simple multilayer perceptrons trained only on per‑sample spike counts can achieve nontrivial accuracy (for example, around 50% for SHD; fig.˜5). This makes it hard to determine whether SNNs truly leverage precise timing or simply exploit rate‑based cues.
Figure 3: Example spike rasters of auditory datasets SHD (top) and SSC (bottom) across the Whole $\rightarrow$ Part $\rightarrow$ Norm normalization stages. Neuron selection and spike count normalization progressively remove rate-based information while retaining temporal structure. Whole (left) is the original dataset. Part (middle) removes neurons whose spike counts are sometimes very low. Norm (right) randomly selects a fixed number of spikes (different for each neuron) leaving only spike timing information and no rate information.
@@
3.2.1 Neuron and Sample Filtering for Spike Count Normalization (Whole $\rightarrow$ Part)
To eliminate rate-based information while preserving temporal structure, we aim to normalize the spike count for each neuron across all samples. However, some neurons emit zero spikes in certain samples, making exact normalization infeasible. In this first stage, we therefore remove those neurons where the minimum spike count is too low.
@@ 3.2.1 Neuron and Sample Filtering for Spike Count Normalization (Whole $\rightarrow$ Part)
Neurons with $c_{i}^{\min}<\theta$ (we use $\theta=2$) are initially marked for exclusion. However, in many cases, a neuron’s low $c_{i}^{\min}$ is caused by just a small number of inactive samples. To retain more neurons, we apply an additional sample filtering step: if a neuron’s low activity affects fewer than a threshold fraction of samples, we remove those samples instead of the neuron. Formally, for each neuron $i$, we define the set of affected samples:
@@
$\text{if }\frac{|\mathcal{M}_{i}|}{M}<\epsilon,\text{ then exclude all samples in }\mathcal{M}_{i}.$ (5)
@@
After filtering, we retain only neurons that spike at least $\theta$ times in every remaining sample, forming the Part variant. To ensure a fair comparison across classes, we randomly downsample each class to match the size of the smallest one. This balancing step improves the interpretability of accuracy metrics in the subsequent experiments.
3.2.2 Min-Count Spike Normalization (Part $\rightarrow$ Norm)
@@
To fully eliminate spike count information while preserving spike timing, we subsample each spike train to retain exactly $c_{i}^{\prime\min}$ spikes per neuron $i$ in every sample, where $c_{i}^{\prime\min}$ is the minimum spike count across the filtered samples defined in the previous stage. Specifically, we randomly select $c_{i}^{\prime\min}$ spike times from each neuron’s spike train in each sample, ensuring fixed per-neuron spike counts across the dataset. This process removes all rate-based information while retaining precise spike timing within each retained spike. However, it may distort global temporal patterns such as inter-spike intervals (ISIs) or synchrony between neurons, depending on which spikes are retained. The resulting Norm dataset preserves the shape of the Part variant and serves as a strict test of spike timing-based learning.
@@
3.3 Surrogate GD Enables SNNs to Exploit Spike Timing in Realistic Datasets
@@
Figure 4: Comparison of SNN architectures trained with surrogate gradient descent (SGD) and SGD-delay (with delays). The top row shows a model with no delay components. The bottom row inserts learnable axonal delays between layers.
@@
Figure 5: Test accuracy under spike timing perturbation $f$ on SHD (left) and SSC (right) spiking speech datasets for models with (solid lines) and without delays (dashed lines). Performance is shown for the original dataset (whole=blue), on a subset of neurons selected for having a minimum number of spikes (part=green), and on a variant where spike count information has been fully removed (norm=red). For the norm variant, spike count information is fully removed, so the accuracy of an MLP trained on spikes counts (grey dashed) corresponds to chance level (5% for SHD, 2.8% for SSC).
We evaluate SGD and SGD-delay on each variant (whole/part/norm), on both the SHD and SSC datasets, across the full range of spike timing perturbations $f\in[0,1]$ (fig.˜5). In every case where spike timing information has not been fully removed by perturbation ($f<1$), the spiking neural network model outperforms the MLP trained on the spike counts (which in turn give chance level performance for the Norm variant of each dataset). As $f\rightarrow 1$ the performance of the SNNs approaches the performance of the MLP using firing rates only. This confirms that surrogate gradient based-learning is not limited to learn rate-based features but can indeed learn from precise spike timing, both with and without explicit delay modeling.
Notably, SGD-delay consistently achieves higher accuracy than SGD across both datasets. For instance, on SHD-norm at $f=0$, SGD achieves 23% while SGD-delay reaches 48%; on SSC-norm, the scores are 15% and 33% respectively. Since we have already seen that delays do not substantially improve the ability of SNNs to extract ISI or synchrony, these results suggest that both SHD and SSC datasets contain substantial long cross-channel ISI information. However, we cannot rule out the possibility that the timing information is in another form that we have not considered but that SNNs are able to learn. These improvements with delays come at a very low parameter cost. The SGD model has $n_{i}n_{h}+n_{h}^{2}+n_{h}n_{o}$ parameters (where $n_{i}$ is the input size which depends on the dataset and variant, $n_{h}=128$ is the hidden layer size, and $n_{o}$ is the number of classes which depends on the dataset). The delays in SGD-delay add only $2n_{h}$ additional parameters. In our case, we increase the number of parameters by at most 0.5%.
@@
Across both SHD and SSC datasets, we observe a common trend in SGD-delay’s performance: as $f$ increases, the degradation in accuracy accelerates sharply—suggesting strong reliance on high-fidelity spike timing. Around f=0.8f{=}0.8, the curves rapidly drop to the corresponding MLP baseline, after which the decay flattens out. By contrast, SGD performance without delays has a more steady decline with increasing $f$ suggesting it uses both fewer and less robust spike timing cues.
4 SNNs With Delays Are More Sensitive to Cross-Channel and Temporal Order Cues
We have seen that networks using delays seem to be better able to make use of information presented at a range of timescales and involving temporal order / causality (section˜2.3), and that this appears to manifest in better performance in the SHD/SSC datasets (section˜3.3). We now probe these insights further with a series of perturbations designed to assess which cues are being used by these models.
4.1 Models
All experiments use the same SNN architecture, training parameters, and loss functions as described in Section 3. For the jitter and deletion analyses below, new models are trained from scratch under each perturbation setting, motivated by the fact that jitter and deletion are likely present in biological systems at training time. In contrast, the time-reversal analysis below reuses the pretrained networks from Section 3, performing inference on temporally inverted versions of the original datasets, to test out-of-distribution performance. Both model variants (SGD and SGD-delay) are evaluated on the SHD and SSC datasets following the same experimental protocol.
4.2 Biologically Inspired Perturbations
To further evaluate the temporal sensitivity of surrogate gradient-trained SNNs, we move beyond the random spike replacement used in earlier sections.
First, we look at per-spike jitter, adding a zero-mean Gaussian noise to each input spike time (clipping to the stimulus duration to avoid spike loss). This type of jitter is widely observed in biological neural networks, and we vary the standard deviation $\sigma$ from 0 to 25 ms.
Second, in order to probe the difference between single-neuron timing cues (like ISI) and cross-neuron cues (like CCISI), we also implemented a per-neuron jitter, identical to the per-spike jitter except that the jitter is re-used for all spikes from a given neuron. This type of jitter is less biologically plausible, but leaves single-neuron timing cues unchanged.
Finally, since many spikes fail in vivo we study performance at different levels of per-spike deletion, where each spike is independently deleted with probability $p_{d}$. While the first two types of jitter affect timing but leave spike rate unchanged, this introduces noise for both timing and rate.
@@ 4.2 Biologically Inspired Perturbations
Figure 6: Spiking neural network performance under temporal perturbations: (a) Per-spike jitter, where each spike is perturbed independently, (b) per-neuron jitter, where all spikes from the same neuron are perturbed equally (c) per-spike deletion. Top: spike rasters of one specific SHD-norm sample before (black) and after (red) perturbation. Bottom: test accuracy across three dataset variants (Whole, Part, Norm), using SGD with and without delay learning.
4.3 Delay-Based Networks Use More Cross-Channel Timing Information
We evaluate model robustness under per-spike jitter, per-neuron jitter and spike deletion on all three SHD variants: whole, part, and norm (fig.˜6). First, we compare per-spike versus per-neuron jitter at a range of noise levels ($\sigma,\sigma_{N}\leq 25$ ms). For per-spike jitter, the delay-based model is able to achieve consistently higher performance than the spike rate-only MLP baseline for all noise levels, while the network without delays falls to or below the MLP level (fig.˜6(a)). By contrast, for per-neuron jitter, performance falls to the MLP level both with and without delays (fig.˜6(c)). Notably, the delay-based model appears much more highly disrupted for the per-neuron jitter than the per-spike jitter. This suggests that the delay-based network is making more use of cross-channel timing information. This is consistent with the differences we saw earlier (section˜2.3), although here the total change to the timing is much smaller. Performance drops significantly by $\sigma_{N}=10$ ms whereas we only saw big differences for the CCISI task when the maximum ISI was larger than 200 ms. Although not directly comparable, this may suggest that temporal order over small time intervals is a particularly important cue for delay-based networks solving auditory tasks.
Finally, we see that both model variants are fairly robust to spike deletion (fig.˜6(c)), suggesting no particular difference in their overall robustness to noise.
4.4 Delay-Based Networks Behave More Like Humans Under Time Reversal
@@
Humans are highly sensitive to the causal structure of sensory input—e.g., speech played backward becomes unintelligible (Saberi and Perrott, 1999). To examine whether SNNs exhibit similar temporal sensitivity, we conduct a time reversal experiment by reversing spike trains within their active time windows (fig.˜7(a)).
@@
$t_{\mathrm{start}},\ t_{\mathrm{end}}]=\left[\min\{t\ |\ x_{i,t}=1\},\ \max\{t\ |\ x_{i,t}=1\}\right],$ (6)
@@
Figure 7: Test accuracy under spike timing perturbation $f$ for original and time-reversed SHD variants. (a) Spike raster of a single sample before and after temporal inversion. (b) Results for models trained without delays. (c) Results for models trained with delays. Each curve corresponds to one of the whole, part, or norm datasets in both original and inverted forms. The MLP level is the performance of a multilayer perceptron trained on spike counts only (no temporal information).
We train on the original data sets and test on the time-reversed versions. This makes sense when comparing to humans, since we do not listen to a lot of time reversed speech (even if we are big Twin Peaks fans). It also leaves all single neuron and synchrony timing cues unchanged, only affecting temporal order / causal cues, and therefore allows us to assess the extent to which the models make use of these different types of temporal cues. We also test at a variety of levels of perturbation $f$ (as in sections˜2 and 3).
Our first finding is that the models without delays are much less affected by time reversal than the models trained with delays, across all variants (fig.˜7(b-c)). This may suggest that delay-based models are a better model of human auditory processing than models without delays.
Our second finding is that for both models, there is a consistent difference between the whole and norm dataset variants. In both cases, under time reversal and no spike timing perturbation ($f=0$), performance is worse than the rate-only MLP baseline. By contrast, for the norm variant, performance still drops but remains better than the MLP baseline. In every case, performance with and without delays matches the MLP performance when all spike timing information is removed ($f=1$). This difference between the whole and norm variants raises a challenge for interpreting these results. We have seen that the whole variant maintains a lot of non-temporal information that is removed by construction in the norm variant. However, it also seems that the whole variant contains much more temporal order structure than the norm variant.
@@ 5 Discussion
In particular, we looked at the Spiking Heidelberg Digits and Spiking Speech Commands datasets (SHD, SSC). These have become widely used benchmarks in the field of spiking neural networks and neuromorphic computing. We find that these datasets do contain substantial and exploitable temporal information, but that it is also possible to perform surprisingly accurate classification using spike counts only, or following various types of perturbation of the spike timing. As a consequence, this dataset may not be ideal if we wish to probe the extent to which various spike-based models can exploit temporal information. We built derived datasets (SHD-norm and SSC-norm) which, by construction, eliminate all spike count information. (Note that although the only information remaining in this dataset is temporal, it is likely that a significant amount of temporal information is also removed by the transformations applied.) We have made a public release of this dataset to facilitate further research on timing-based computation in spiking neural networks.
@@
The code for loading the variant SHD and SSC datasets is available on the GitHub repository: https://github.com/neural-reckoning/temporal-shd. The datasets are archived on Zenodo(Yu et al., 2025) (https://doi.org/10.5281/zenodo.16153275).